OpenStudy (anonymous):

Jackson deposits $160 each month into a savings account earning interest at the rate of 8% per year compounded monthly. How much will he have in this account at the end of 7 years? a) $17,940.13 b) $17,938.13 c) $17,937.13 d) $17,936.13 e) $17,934.13 f) None of the above.

OpenStudy (mathmate):

monthly interest, i = 0.08/12 monthly deposit, D = 160 number of periods(months), n = 7*12 = 84 Final amount, A \[ A = \frac{D(1+i)^n - 1)}{i} \]